OpenStudy (anonymous):

The number of subsets that can be created from the set {1, 2, 3} is 3 6 7 8

jimthompson5910 (jim_thompson5910):

hint: if there are n elements in the set, then there are \(\Large 2^n\) different subsets

OpenStudy (anonymous):

whaaaaat..

jimthompson5910 (jim_thompson5910):

example the set {a, b, c, d, e} has 5 elements, so n = 5 Therefore, there are 2^5 = 32 different subsets

OpenStudy (anonymous):

confused..but thanks anyways

jimthompson5910 (jim_thompson5910):

how many elements are in the given set {1, 2, 3}

OpenStudy (anonymous):

I don't know..

jimthompson5910 (jim_thompson5910):

how many numbers are in that set

OpenStudy (anonymous):

3

jimthompson5910 (jim_thompson5910):

so n = 3

jimthompson5910 (jim_thompson5910):

which means 2^n = 2^3 = 2*2*2 = 8

jimthompson5910 (jim_thompson5910):

So there are 8 different subsets

jimthompson5910 (jim_thompson5910):

And they are... {1, 2, 3} {1, 2}, {1, 3}, {2, 3} {1}, {2}, {3} { } The last set { } is the empty set and it's a subset of any set.

OpenStudy (anonymous):

thank you so much.

jimthompson5910 (jim_thompson5910):

yw